4 Suppose that the civilian non-institutional working age population is 244,863, the civilian labor force 155,654, and the employment to population ratio 58.6%. a)What is the unemployment rate? b What is the labor force participation rate? c) If the civilian non-institutional population grows by 1% over the next year and the labor force participation rate remains constant, how many jobs must be created over the next year to leave the unemployment rate unchanged? d) If there is no change in the civilian non-institutional population over the next year, but the labor force participation rate rises by 1%, how many jobs must be created over the next year to leave the unemployment rate unchanged? 5. Suppose the government were to subsidize the wages of all women in the population by paying their employers 50 cents for every hour the worked. What would be the effect on the wage rate women received? What would be the effect on the net wage employers paid? Show this on a graph. On a separate graph, consider the labor market for men. Label both axes and show the equilibrium wage and employment levels. Would the subsidy for women affect the labor market for men? Why or why not? If so, show this on your graph. 6.
Answer #1

Answer 4 (a) : The Total civilian non-institutional working age population = 244,863

                          The total labor force available = 155,654

                           The employment rate is given as = 58.6% of the total population, which means:

                            The number of individuals employed = 58.6/100 x 244,863

                                                                                               = 143,489

                             Hence 143,489 individuals are employed. Now, these 143,489 must be among the labor force. Therefore, the number of labor force who are not employed is = 155,654 – 143,489

                                                                                                  = 12,165

                             Therefore, the unemployment rate will be = (100 /155,654) x 12165

                                                                                                         = 7.8

Therefore, the unemployment rate is 7.8%

Answer (b) : The Labor force participation rate means the number of labor who are actively participating in the industry. Here, since 7.8 % of the total labor force is unemployed, therefore the rest of the labor force must be part of the actively participating employed group. Therefore, the labor force participation rate is 92.2 %.

Answer ( c) : If the civilian non-institutional population grows by 1% , this means the total civilian non-institutional population will be = 244,863 + ( 1% of 244,863 )

                                                        = 244,863 + 2448

                                                        = 247,311

     If the labor force participation rate remains constant, that means 92.2% , therefore, the labor force employed out of the total population which is 58.6% is = 58.6 % of 247,311

                                                                                                  = (58.6/100) x 247,311

                                                                                                  = 144,924

Therefore, after the population rises by 1%, number of labor force is 144,924.

Now, since 92.2% population is working, that means = 92.2% of 144,924

                                                                                              = 133619

Therefore the remaining of the pollution not working is = 144924 – 133619

                                                                                                    = 11305

   Hence 805 more jobs have to be created in the next year.


Answer (d) : If the labor force participation grows by 1% , this means the labor force participating will be = 155,654 + ( 1% of 155,654 )

                                      = 155654 + 1556

                                      = 157,210

Therefore , 1556 number of jobs must at least be created next to keep the unemployment rate unchanged.